A copper wire has cross sectional area of 6.7 × 10-8 m2
and is 1 m long. Calculate the
current flowing in the wire if the wire is connected to a 9 V battery. The resistivity of
copper is given as 1.68 × 10-8 Ω
1
Expert's answer
2021-09-29T09:55:25-0400

Given:

"A=6.7*10^{-8}\:\rm m^2"

"l=1.0\:\rm m"

"V=9\:\rm V"

"\rho=1.68*10^{-8}\:\rm\Omega\cdot m"


The electric current

"I=\frac{V}{R}=\frac{V}{\rho\frac{l}{A}}"

"I=\frac{9}{1.68*10^{-8}*\frac{1.0}{6.7*10^{-8}}}=35.9\:\rm A"